chn The anetage molecular speed ma sample ofCH. pas at a certan tmperatge s 6 ms The average molecula speed in a sample of NH g
Answer #1

molar mass of CH4 ( M1) = 16g/mole

molar mass of NH3 (M2) = 17g/mole

The average molecular speed of CH4   (C1) = 676m/sec

The average molecular speed of NH3 (C2)   =

C1/C2     = \sqrt{}M2/M1

676/C2    = \sqrt{}17/16

676/C2      = 1.0307

C2            = 676/1.0307   = 655.86m/sec

The average molecular speed of NH3 (C2)   = 655.86m/sec